# Edge States with Hidden Topology in Spinner Lattices

### v1

@@ Abstract
Symmetries—whether explicit, latent, or hidden—are fundamental to understanding topological materials. This work introduces a prototypical spring-mass model that extends beyond established canonical models, revealing topological edge states with distinct profiles at opposite edges. These edge states originate from hidden symmetries that become apparent only in deformation coordinates, as opposed to the conventional displacement coordinates used for bulk-boundary correspondence. Our model realized through the intricate connectivity of a spinner chain, demonstrates experimentally distinct edge states at opposite ends. By extending this framework to two dimensions, we explore the conditions required for such edge waves and their hidden symmetry in deformation coordinates. We also show that these edge states are robust against disorders that respect the hidden symmetry. This research paves the way for advanced material designs with tailored boundary conditions and edge state profiles, offering potential applications in fields such as photonics, acoustics, and mechanical metamaterials.
@@ INTRODUCTION
Recent research has unveiled that symmetries can also manifest in latent [24, 25] or hidden forms [26, 27, 28, 29, 30]. For instance, in spring-mass models—widely applicable across mechanical, optical, acoustic systems, and electrical circuits—a mass dimer chain with an odd number of particles and free boundaries is known to support edge states [31, 32]. However, the topological origin of these states was not well understood. Allein et al. [30] demonstrated that the underlying chiral symmetry for topological characterization emerges in deformation coordinates rather than the traditional displacement coordinates used for spring-mass models. This insight opens new avenues for designing material architectures by considering complex spring connections [33, 34, 35, 36, 37, 38] and varying masses [39, 40], suggesting that modeling dynamics in deformation coordinates could reveal novel topological phases in systems with hidden symmetries.
Canonical spring-mass models, such as stiffness and mass dimers, support topological edge states protected by chiral symmetry, which are revealed in displacement and deformation coordinates, respectively. These models require different particle numbers (even for stiffness dimers and odd for mass dimers) and distinct boundary conditions (fixed-fixed for stiffness dimers and free-free for mass dimers) to exhibit symmetry-protected edge states. Despite their importance, both systems share a common feature: mirror symmetry, which implies identical edge state profiles on opposite edges. Discovering a new prototypical mass-spring model that supports topologically robust edge states at opposite ends with distinct profiles could significantly expand the design space of topological materials.
To address this, we explore more complex spring-mass models. We present a prototypical model incorporating two distinct masses and two types of connectivity, supporting edge states with distinct profiles at free edges. Traditionally, different masses have been used to break mirror symmetry. Here, we show that the bipartite nature of masses can be viewed not only as a mechanism for breaking mirror symmetry but also as a tuning tool for hidden topological effects. We propose a chain of spinners strategically interconnected by springs as a physical realization of this model. We experimentally demonstrate different edge states at opposite ends, with their topological nature revealed through deformation coordinate analysis. We also extend this framework to two dimensions, highlighting the intricate fine-tuning required to observe the hidden symmetry of propagating waves on opposite edges of the lattice. Additionally, we perform a disorder analysis based on the underlying hidden symmetry, showing that symmetry-protected disorder can be leveraged to develop material architectures supporting robust topological edge states.
This work enhances our understanding of hidden symmetries in topological systems and opens new pathways for designing advanced materials with tailored boundary conditions and edge state profiles, potentially impacting fields such as photonics, acoustics, and mechanical metamaterials.
@@
where $u_{1,n}$ and $u_{2,n}$ represent the displacements of the first and second particle in the $n$th unit cell. Upon substituting plane wave solutions, we get the non-dimensional bulk displacement dynamical matrix $\mathcal{D}_{u}(P,q)$ that can be represented in the Pauli basis as follows: ${\mathcal{D}_{u}(P,q)=c_{0}\mathcal{I}+c_{1}\sigma_{1}+c_{2}\sigma_{2}+c_{3}\sigma_{3}}$, where $c_{0}=1+P$, $c_{1}=-(1+P\cos{q})$ ), $c_{2}=-P\sin{q}$ ), ${c_{3}=0}$, $P\coloneqq k_{2}/k_{1}$, and $q\in[-\pi,\pi]$ is the wavenumber. $\mathcal{D}_{u}(P,q)$ possesses chiral symmetry, i.e., ${\sigma_{3}\hat{\mathcal{D}}_{u}(P,q)\sigma_{3}=-\hat{\mathcal{D}}_{u}(P,q)}$, where $\hat{\mathcal{D}}_{u}(P,q)$ is the trace-less counterpart. The dynamical matrix maps to the Hamiltonian of the standard SSH chain with two different hopping strengths and an onsite potential. Consequently, the spectrum is symmetric about a finite frequency, and the eigenvectors $u_{\pm}$ of the acoustic and optical bands form chiral pairs. The system has a well-defined topological invariant, namely, a winding number. A nonzero winding for $P>1$ indicates non-trivial topology and a corresponding finite lattice with even particles and fixed edges preserves chiral symmetry and supports boundary states on the left and right ends. Since the finite lattice preserves the mirror symmetry, both boundary states also have the same localization profiles.
@@
$\sigma_{1}\mathcal{D}_{e}[P,\beta,\boldsymbol{q}=(q_{x},q_{y})]\sigma_{1}=\mathcal{D}_{e}[P,\beta,\boldsymbol{q}=(-q_{x},q_{y})].$
@@
In Fig. 3D, we plot the spectrum of a ribbon consisting of 41 spinners, with free boundaries along the $x$-axis and Floquet boundary conditions along the $y$-axis (see the 2D localization parameter in Materials and Methods). We observe localized states on the left and right boundaries of the ribbon in the band gap for all values of $q_{y}\in[-\pi,\pi]$. We then consider a finite structure with 31 and 16 spinners along the $x$ and $y$-axes, respectively. In Fig. 3E, we plot the spectrum of the finite 2D system with free (fixed) left-right (top-right) boundaries. We observe multiple edge states in the band gap. These edge states exhibit different localization profiles at the left and right edges, as shown in Fig. 3F, which is a hallmark of our mixed spinner lattice. Upon transformation to deformation coordinates, the chiral-like symmetry is revealed, as shown in Fig. 3G, where every alternate particle is at rest, similar to the 1D chain in the deformation framework. The lattice also reveals the mirror symmetry along the $x$-axis in the deformation framework, similar to the 1D chain.
@@
Looking ahead, the design space provided by our spinner-based models offers a valuable framework for exploring various tight-binding 2D models. Additionally, this approach opens opportunities for investigating novel classes of higher-order topological lattices [36], with diverse boundary conditions and localization profiles, based on hidden symmetries.
MATERIALS AND METHODS

### v2

@@ Abstract
Symmetries—whether explicit, latent, or hidden—are fundamental to understanding topological materials. This work introduces a prototypical spring-mass model that extends beyond established canonical models, revealing topological edge states with distinct profiles at opposite edges. These edge states originate from hidden symmetries that become apparent only in deformation coordinates, as opposed to the conventional displacement coordinates used for bulk-boundary correspondence. Our model, realized through the intricate connectivity of a spinner chain, demonstrates experimentally distinct edge states at opposite ends. By extending this framework to two dimensions, we explore the conditions required for such edge waves and their hidden symmetry in deformation coordinates. We also show that these edge states are robust against disorders that respect the hidden symmetry. This research paves the way for advanced material designs with tailored boundary conditions and edge state profiles, offering potential applications in fields such as photonics, acoustics, and mechanical metamaterials.
@@ INTRODUCTION
Recent research has unveiled that symmetries can also manifest in latent [24, 25] or hidden forms [26, 27, 28, 29, 30]. For instance, in spring-mass models—widely applicable across mechanical, optical, acoustic systems, and electrical circuits—a mass dimer chain with an odd number of two altering masses and uniform stiffness, and with free boundaries is known to support edge states [31, 32]. However, the topological origin of these states was not well understood. Allein et al. [30] demonstrated that the underlying chiral symmetry for topological characterization emerges in deformation coordinates rather than the traditional displacement coordinates used for spring-mass models. This insight opens new avenues for designing architectured materials by considering varying masses which could support novel topological states protected by hidden chiral symmetry.
In this work, we enlarge the design space of topological lattices by varying mass along with different types of spring connections within the unit cell. We show that by using complex spring connections [33, 34, 35, 36, 37, 38], mass dimer 1D and 2D lattices could support distinct and unconventional profiles of topologically robust edge states at opposite ends—a phenomenon not observed in canonical two-band models [39, 40, 41, 30, 42]. Such profiles result from the breakage of both chiral and mirror symmetries. However, we show that the edge states are topological since both chiral and mirror symmetries are revealed in deformation coordinates.
Furthermore, we propose a chain of spinners strategically interconnected by springs as a physical realization of this model. We experimentally demonstrate different edge states at opposite ends, with their topological nature revealed through deformation coordinate analysis. We also extend this framework to two dimensions, highlighting the intricate fine-tuning required to observe the hidden symmetry of propagating waves on opposite edges of the lattice. Additionally, we perform a disorder analysis based on the underlying hidden symmetry, showing that symmetry-protected disorder can be leveraged to develop material architectures supporting robust topological edge states.
This work enhances our understanding of hidden symmetries in topological systems and opens new pathways for designing advanced materials with tailored edge state profiles, potentially impacting fields such as photonics, acoustics, and mechanical metamaterials.
@@
where $u_{1,n}$ and $u_{2,n}$ represent the displacements of the first and second particle in the $n$th unit cell. Upon substituting plane wave solutions, we get the non-dimensional bulk displacement dynamical matrix $\mathcal{D}_{u}(P,q)$ that can be represented in the Pauli basis as follows: ${\mathcal{D}_{u}(P,q)=c_{0}\mathcal{I}+c_{1}\sigma_{1}+c_{2}\sigma_{2}+c_{3}\sigma_{3}}$, where $c_{0}=1+P$, $c_{1}=-(1+P\cos{q})$ ), $c_{2}=-P\sin{q}$ ), ${c_{3}=0}$, $P\coloneqq k_{2}/k_{1}$, and $q\in[-\pi,\pi]$ is the wavenumber. $\mathcal{D}_{u}(P,q)$ possesses chiral symmetry, i.e., ${\sigma_{3}\hat{\mathcal{D}}_{u}(P,q)\sigma_{3}=-\hat{\mathcal{D}}_{u}(P,q)}$, where $\hat{\mathcal{D}}_{u}(P,q)$ is the trace-less counterpart. The dynamical matrix maps to the Hamiltonian of the standard SSH chain with two different hopping strengths and an onsite potential. Consequently, the spectrum is symmetric about a finite frequency, and the eigenvectors $u_{\pm}$ of the acoustic and optical bands form chiral pairs. The system has a well-defined topological invariant, namely, a winding number. A nonzero winding for $P>1$ indicates non-trivial topology, and a corresponding finite lattice with even particles and fixed edges preserves chiral symmetry and supports boundary states on the left and right ends. Since the finite lattice preserves the mirror symmetry, both boundary states also have the same localization profiles.
@@
$\sigma_{1}\mathcal{D}_{e}\left(P,\beta,\boldsymbol{q}=(q_{x},q_{y})\right)\sigma_{1}=\mathcal{D}_{e}\left(P,\beta,\boldsymbol{q}=(-q_{x},q_{y})\right).$
@@
In Fig. 3D, we plot the spectrum of a ribbon consisting of 41 spinners, with free boundaries along the $x$-axis and Floquet boundary conditions along the $y$-axis (see the 2D localization parameter in Materials and Methods). We observe localized states on the left and right boundaries of the ribbon in the band gap for all values of $q_{y}\in[-\pi,\pi]$. We then consider a finite structure with 31 and 16 spinners along the $x$ and $y$-axes, respectively. In Fig. 3E, we plot the spectrum of the finite 2D system with free (fixed) left-right (top-bottom) boundaries. We observe multiple edge states in the band gap. These edge states exhibit different localization profiles at the left and right edges, as shown in Fig. 3F, which is a hallmark of our mixed spinner lattice. Upon transformation to deformation coordinates, the chiral-like symmetry is revealed, as shown in Fig. 3G, where every alternate particle is at rest, similar to the 1D chain in the deformation framework. The lattice also reveals the mirror symmetry along the $x$-axis in the deformation framework, similar to the 1D chain.
@@
Looking ahead, the design space provided by our spinner-based models offers a valuable framework for exploring various tight-binding 2D models. Additionally, this approach opens opportunities for investigating novel classes of higher-order topological lattices [36], with diverse boundary conditions and localization profiles based on hidden symmetries.
METHODS
@@
Author contributions – U.V. and R.C. conceived the project. U.V., under the supervision of G.T. and R.C., conceptualized the spinner-based models and developed the theoretical framework. M.I. and U.V. fabricated the setup and conducted the experiments. U.V. analyzed the experimental data. U.V., G.T., and R.C. contributed to the manuscript writing. R.C. supervised the overall project.